OpenStudy (anonymous):

Cassie has 30 grams of cobalt-60 that decays 12% per year. How many years will it take for half of the original amount to decay??? * PLEASE HELP :)*

OpenStudy (lena772):

100-12=88% Original amount=30 Half of this=15 1st year 30*88%=26.4 2nd year 26.4*88%=23.232 3rd year 23.232*88%=20.44416 4th year 20.44416*88%=17.9908608 5th year 17.99086087*88%=15.8319575656 6th year 15.8319575656*88%=13.9321226577 A little over 5 years, sometime during the 6th year, half of the original amount will be decayed.

OpenStudy (aaronq):

You can find the exact amount of time it will take with those equations: \(ln[A]_t=-kt+ln[A]_o\) and \(t_{1/2}=\dfrac{ln(2)}{k}\) Find the half life of the substance.